# Prompt-prompted Mixture of Experts for Efficient LLM Generation

### v2

@@ Abstract
With the development of transformer-based large language models (LLMs), they have been applied to many fields due to their remarkable utility, but this comes at a considerable computational cost at deployment. Fortunately, some methods such as pruning or constructing a mixture of experts (MoE) aim at exploiting sparsity in transformer feedforward (FF) blocks to gain boosts in speed and reduction in memory requirements. However, these techniques can be very costly and inflexible in practice, as they often require training or are restricted to specific types of architectures. To address this, we introduce GRIFFIN, a novel training-free MoE that selects unique FF experts at the sequence level for efficient generation across a plethora of LLMs with different non-ReLU activation functions. This is possible due to a critical observation that many trained LLMs naturally produce highly structured FF activation patterns within a sequence, which we call flocking. Despite our method’s simplicity, we show with 50% of the FF parameters, GRIFFIN maintains the original model’s performance with little to no degradation on a variety of classification and generation tasks, all while improving latency (e.g. 1.25$\times$ speed-up in Llama 2 13B on an NVIDIA L40). Code is available at https://github.com/hdong920/GRIFFIN.
@@ 1 Introduction
Transformers [${}^{+}$17] have demonstrated incredible capabilities across a plethora of domains [LWLQ22, ${}^{+}$22, ${}^{+}$23]. Their large language model (LLM) successors [${}^{+}$23, ${}^{+}$23, ${}^{+}$23, ${}^{+}$24, ${}^{+}$24, Ant24] have pushed the bar higher, but these behemoths have become performative at the price of enormous amounts of computation and memory demands. One significant contributor is the model size itself. Not only is storage an issue, model layers tend to be wide and plenty, slowing down inference. Moreover, given the existence of sparse structures in LLMs, especially in feedforward (FF) blocks [GSBL20, DLBZ22, ${}^{+}$22, ${}^{+}$23], these models waste computation on intermediate features with little to no impact on the final result. For instance, it has been observed that in OPT-175B [${}^{+}$22], fewer than 5% of neurons in FF blocks have nonzero values per token [${}^{+}$23], meaning 95% of the compute in each FF block is wasted. Usually consisting of around two-thirds of the parameters in an LLM, FF blocks can be serious memory and compute bottlenecks. These inefficiencies are highly problematic in latency-sensitive scenarios like in chatbots and autonomous vehicles.
There have been many methods to exploit sparsity in LLMs for efficiency gains, such as pruning model weights and constructing mixtures of experts (MoEs). Pruning removes low-impact pre-trained weights to reduce storage, yet this often does not translate to real speed-ups in practice, unless the pruning is done in a hardware-friendly manner which typically causes greater performance deterioration. MoEs better preserve the original performance by adaptively selecting subsets of the model to use per input but also come with drawbacks. Unless the model has been trained in this fashion [FZS22, ${}^{+}$24], it will need to learn a cheap yet effective gating function (expert selection mechanism) and sometimes even require full fine tuning. Perhaps an even bigger weakness of many of these methods is the inability to effectively carry over to pre-trained LLMs with non-ReLU activations. We seek to overcome these challenges with MoEs:
@@
    Most current LLMs are not MoEs, so they must be adapted by training gating functions and/or fine tuning the whole model.
@@
    Gating functions should be simple enough to not cost more than they save.
@@
    Many current approaches work solely on FF blocks with ReLUs or depend on expensive conversions to ReLUs.
@@
Figure 1: Relative FF activation magnitudes of the first 512 features and tokens across a sequence from PG-19 [${}^{+}$19, ${}^{+}$20] in layer 10 of Llama 2 7B (left) and Gemma 7B (right). These heatmaps show flocking, where relative activation magnitudes are shared within a sequence. More examples in Appendix B.
Unlike existing pruning or MoE methods, we exploit flocking in our design of GRIFFIN (Gating by Repetition In Feedforward Intermediate Neurons), a highly performative and efficient training-free method to turn an LLM into an MoE. GRIFFIN does this by using a sequence’s prompt to determine the experts to activate during generation, allowing it to overcome all of the aforementioned challenges:
@@
    No Preparation: Our no-cost method is completely training-free and requires no preparation. Moreover, the simple implementation of GRIFFIN means it can easily be dropped into any FF block.
@@
    Simple Expert Selection: Since flocking exists throughout a sequence, the prompt reveals the most relevant FF neurons for generation with little to no performance loss. The selection process is parameter-free and adds negligible overhead.
@@
    Model & Activation Function Diversity: Thorough experimentation demonstrates the efficacy of GRIFFIN on numerous models, including Llama 2 [${}^{+}$23], Gemma [${}^{+}$24], Mistral [${}^{+}$23], OPT [${}^{+}$22], and ReluLlama [Tea23]. Together, the tested activation functions include ReLU, SwiGLU, GEGLU, and ReGLU [Sha20].
In this paper, we show GRIFFIN is a simple and strong method to turn any LLM into MoE because of flocking. In the next section (Section 2), we discuss some strengths and weaknesses of current methods that seek to improve FF efficiency. In more detail, we formalize the MoE problem and its motivation in Section 3 and present our novel approach in Section 4.2, which requires a thorough examination of the surprising phenomenon of flocking shared by many LLMs in Section 4.1. Our rigorous experiments demonstrate GRIFFIN preserves performance on classification and generation even after removing 50% of FF neurons (Section 5.1), all while having lower latency (Section 5.2). For instance, GRIFFIN reduces the number of active parameters in Llama 2 13B from 13B to 8.8B during generation to improve latency by 1.25$\times$ with almost no loss in performance. Finally, we show our method’s incredible scalability and robustness in several ablation studies (Section 5.3).
@@ 2 Background
The observation that transformer FF blocks produce sparse activations is not new [GSBL20, DLBZ22, ${}^{+}$22, DCC23, ${}^{+}$23]. In ReLU-based LLMs like OPT [${}^{+}$22], the activations can be exceptionally sparse and become more apparent for larger models [${}^{+}$23]. As more models use non-sparse activation functions like GLU variants [Sha20], it is difficult for neurons to have no contribution to the output since these functions do not have an interval that maps to zero. Without exact sparsity, the efficacy of these methods becomes limited. As such, this has ushered a wave of models that are either adapted from available models [${}^{+}$21, LLLC21, ${}^{+}$23, ${}^{+}$24, ${}^{+}$24] or trained from scratch [FZS22] which can produce activations that are exactly zero with little to no performance loss. Even so, these methods require considerable amounts of computational resources.
@@
Pruning [LDS89] is another sparsity-guided way to tackle compute and memory bottlenecks of models. Previously, the common method would be some variation of iteratively rounding weights down to zero based on some score and retraining to recover any lost performance [FC18, BGOFG20, ${}^{+}$21, ${}^{+}$24]. While this can result in most parameters being pruned, this method comes with a few issues. First, with the increasing scale of LLMs, retraining becomes impractical for most. Fortunately, cheap methods to effectively prune LLMs have been developed [FA23, SLBK23, ${}^{+}$23, ${}^{+}$24]. The second issue is that unless pruning is done in a structured manner [XZC22, SWSL23, MFW23, ${}^{+}$23, XGZC23], it is difficult to see real computational savings, yet structured pruning often leads to much more severe performance degradation. Third, pruning enforces sparsity to be static which can be strong assumption since FF blocks are widely believed to contain the model’s memory [GSBL20].
@@
Making sparsity more dynamic has motivated the design of mixture-of-experts (MoEs) [JJNH91] to avoid computing low-impact features in trained models at varying granularities [${}^{+}$21, ${}^{+}$23, PSBW23, CIS23, ${}^{+}$23, ${}^{+}$24, ${}^{+}$24]. The main idea of MoEs is to use a gating function to identify a small subset of neurons that will be used to compute the layer’s output, an adaptive form of pruning. In the ideal case, all active neurons are selected and inactive neurons are ignored for each input. However, current methods either require training or rely on ReLU activation functions. Our method has the best of both worlds: it is training-free and effective on a variety of activation functions.
@@ 3 Problem Formulation
This section contains an overview of different components of the FF block followed by a more detailed introduction of the MoE problem which our method aims to tackle. Since FF blocks operate identically and independently for each token unlike attention, we begin with defining the FF block with a single column vector input $\bm{x}\in\mathbb{R}^{D}$:
@@
where $\odot$ signifies element-wise multiplication. For all examples, $\bm{W}_{1},\bm{W}_{g}\in\mathbb{R}^{D_{\text{FF}}\times D}$ and $\bm{W}_{2}\in\mathbb{R}^{D\times D_{\text{FF}}}$ where typically, $D_{\text{FF}}\gg D$. We refer to $\bm{z}=\text{FF}_{1}(\bm{x})$ as the FF activations. The goal of MoEs is to find $\widehat{\bm{W}}_{1}\in\mathbb{R}^{k\times D}$, $\widehat{\bm{b}}_{1}\in\mathbb{R}^{k}$, and $\widehat{\bm{W}}_{2}\in\mathbb{R}^{D\times k}$ (additionally $\widehat{\bm{W}}_{g}\in\mathbb{R}^{k\times D}$ and $\widehat{\bm{b}}_{g}\in\mathbb{R}^{k}$ if needed) where $k<D_{\text{FF}}$ such that when the FF block is reparameterized with these matrices, the output value is preserved. In other words, for
@@
$\text{FF}(\bm{x})\approx\widehat{\text{FF}}_{2}(\widehat{\text{FF}}_{1}(\bm{x}))$, and similarly for FF blocks with non-GLU functions. In the MoE setting, these smaller matrices can vary from token to token and are actually selections of rows and columns of the original structures. Crucially, this selection leads to multiplication with smaller matrices which are naturally efficient on GPUs and TPUs [FSH04, WWB19]. For all equations defined in this section, they operate independently on each row of a length $S$ sequence input $\bm{X}\in\mathbb{R}^{S\times D}$ (e.g. the activations for a sequence are $\bm{Z}=\text{FF}_{1}(\bm{X})\in\mathbb{R}^{S\times D_{\text{FF}}}$).
@@ 4.1 Observing Flocking
Flocking arises when we look at the relative impact of each neuron per token within a sequence. To see this, we normalize rows of $\bm{Z}$ to be unit vectors to construct $\overline{\bm{Z}}\in\mathbb{R}^{S\times D_{\text{FF}}}$ (i.e. $[\overline{\bm{Z}}]_{i}=[\bm{Z}]_{i}/\|[\bm{Z}]_{i}\|_{2}$), the relative activations. We show example relative activation magnitudes for a sequence in Llama 2 7B and Gemma 7B in Figure 1. Since there are distinct vertical streaks, this intriguingly implies that activations that have relatively greater weight are common across all tokens in a sequence. Notably, Llama 2 7B and Gemma 7B use SwiGLU and GEGLU activations, respectively, along with other major architecture differences. We call this phenomenon flocking, like highly organized groups of birds, and we observe this in virtually all FF layers (see Appendix B).
While relative activations magnitudes are shared within a sequence, they are not generally shared between sequences. We show this by taking the $\ell_{2}$-norm of $\overline{\bm{Z}}$ along the token axis to obtain a length $D_{\text{FF}}$ vector for each sample or sequence, roughly capturing the contribution of each FF neuron throughout a sequence. Taking the top-$k$ of this for each sample at each layer, we find the Jaccard similarity between two sequences based on the indices selected for different $k$. In other words, we compute the intersection over union of every unique pair of top-$k$ sets. Higher values indicate more similar top-$k$ sets. From Figure 2 where we aggregate Jaccard similarities across WikiText [MXBS16] samples, we observe a lack of inter-sample activation similarities for the vast majority of layers in Llama 2 7B and Gemma 7B, unless the sets of selected neurons are large. This lack of consistency implies pruning entire FF neurons without retraining would be less effective than a more adaptive method.
@@ 4.2 GRIFFIN Algorithm
Using our insight on flocking, we introduce GRIFFIN as a simple general purpose and training-free MoE method for efficient generation, captured in Figure 3. In a nutshell, we select experts during the prompt phase of each sample which are then used for the entire duration of the generation phase. This effective approach is based on a key observation on flocking: since tokens within a sequence share activation patterns, the prompt and generated tokens will also share activation patterns.
Prompt Phase Expert Selection.
Our experts or neurons are chosen at the sequence level, so we need to consider the dynamics of the entire input sequence rather than just a single token when choosing our experts. To select expert neurons, we need a statistic $\bm{s}\in\mathbb{R}^{D_{\text{FF}}}$ to inform us of the importance of each neuron. At the prompt phase, we do this by aggregating taking the $\ell_{2}$-norm of $\overline{\bm{Z}}$ along the token axis:
@@
Taking the indices of the top-$k$ across $\bm{s}$ gives us the neurons we will use for this sample’s generation phase which make up the set $\mathcal{E}$. Using the experts in $\mathcal{E}$, we can find $\widehat{\bm{W}}_{1},\widehat{\bm{b}}_{1},\widehat{\bm{W}}_{g},\widehat{\bm{b}}_{g},\text{ and }\widehat{\bm{W}}_{2}$ by selecting corresponding rows and columns in $\bm{W}_{1},\bm{b}_{1},\bm{W}_{g},\bm{b}_{g}$, and $\bm{W}_{2}$, respectively. This is done for every FF block during the prompt phase. Elaborated in Appendix A, $\bm{s}$ highlights neurons consistently activated at relatively high intensities.
Generation with Experts.

### v3

# Prompt-prompted Adaptive Structured Pruning for Efficient LLM Generation

@@ Abstract
With the development of transformer-based large language models (LLMs), they have been applied to many fields due to their remarkable utility, but this comes at a considerable computational cost at deployment. Fortunately, some methods such as pruning or constructing a mixture of experts (MoE) aim at exploiting sparsity in transformer feedforward (FF) blocks to gain boosts in speed and reduction in memory requirements. However, these techniques can be very costly and inflexible in practice, as they often require training or are restricted to specific types of architectures. To address this, we introduce GRIFFIN, a novel training-free and calibration-free method that selects unique FF experts at the sequence level for efficient generation across a plethora of LLMs with different non-ReLU activation functions. This is possible due to a critical observation that many trained LLMs naturally produce highly structured FF activation patterns within a sequence, which we call flocking. Despite our method’s simplicity, we show with 50% of the FF parameters, GRIFFIN maintains the original model’s performance with little to no degradation on a variety of classification and generation tasks, all while improving latency (e.g. 1.29$\times$ and 1.25$\times$ speed-ups in Gemma 7B and Llama 2 13B, respectively, on an NVIDIA L40). Code is available at https://github.com/hdong920/GRIFFIN.
@@ 1 Introduction
Transformers [VSP+17] have demonstrated incredible capabilities across a plethora of domains [LWLQ22, KNH+22, NBZ+23]. Their large language model (LLM) successors [TMS+23, TAB+23, JSM+23, JSR+24, TMH+24, Ant24] have pushed the bar higher, but these behemoths have become performative at the price of enormous amounts of computation and memory demands. One significant contributor is the model size itself. Not only is storage an issue, model layers tend to be wide and plenty, slowing down inference. Moreover, given the existence of sparse structures in LLMs, especially in feedforward (FF) blocks [GSBL20, DLBZ22, LYB+22, LWD+23], these models waste computation on intermediate features with little to no impact on the final result. For instance, it has been observed that in OPT-175B [ZRG+22], fewer than 5% of neurons in FF blocks have nonzero values per token [LWD+23], meaning 95% of the compute in each FF block is wasted. Usually consisting of around two-thirds of the parameters in an LLM, FF blocks can be serious memory and compute bottlenecks. These inefficiencies are highly problematic in latency-sensitive scenarios like in chatbots and autonomous vehicles.
There have been many methods to exploit sparsity in LLMs for efficiency gains, such as pruning model weights and constructing mixtures of experts (MoEs). Pruning removes low-impact pre-trained weights to reduce storage, yet this often does not translate to real speed-ups in practice, unless the pruning is done in a structured and hardware-friendly manner [XZC22, SWSL23, MFW23, LYZ+23, XGZC23] which typically causes greater performance deterioration. MoEs better preserve the original performance by adaptively selecting subsets of the model to use per input but also come with drawbacks. Unless the model has been trained in this fashion [FZS22, JSR+24], it will need to learn a cheap yet effective gating function (expert selection mechanism) and sometimes even require full fine tuning. Perhaps an even bigger weakness of many of these methods is the limited effectiveness and special considerations to strictly enforce ReLU-like activation functions [ZLL+21, LYB+22, CIS23]. In summary, pruning and MoEs provide enormous benefits but also come with steep challenges:
@@
    Structured sparsity is difficult and costly to enforce without serious performance degradation.
@@
    Adaptive selection of model subsets need to be cheap and accurate.
@@
    The method should be effective with various activation functions.
@@
Figure 1: Relative FF activation magnitudes of the first 512 features and tokens across a sequence from PG-19 [RPJ+19, GBB+20] in layer 10 of Llama 2 7B (left) and Gemma 7B (right). These heatmaps show flocking, where relative activation magnitudes are shared within a sequence, illustrated with the distinct dark vertical streaks. More examples in Appendix E.
Unlike existing pruning or MoE methods, we exploit flocking in our design of GRIFFIN (Gating by Repetition In Feedforward Intermediate Neurons), a highly performative and efficient training-free method to adaptively activate FF neurons. GRIFFIN does this by using a sequence’s prompt to determine the experts to activate during generation, allowing it to overcome all of the aforementioned challenges:
@@
    No Preparation: Our no-cost method is completely training-free and requires no calibration. Moreover, the simple implementation of GRIFFIN means it can be dropped into any FF block and instantly deployed.
@@
    Simple Expert Selection: Flocking in the prompt reveals the most relevant FF neurons for generation with little to no performance loss. The selection process is parameter-free and adds negligible overhead.
@@
    Model & Activation Function Diversity: Thorough experimentation demonstrates the efficacy of GRIFFIN on numerous models, including Llama 2 [TMS+23], Gemma [TMH+24], Mistral [JSM+23], OPT [ZRG+22], and ReluLlama [Tea23]. Together, the tested activation functions include ReLU, SwiGLU, GEGLU, and ReGLU [Sha20].
In this paper, we show GRIFFIN is a simple and strong adaptive structured pruning method because of flocking. In the next section (Section 2), we discuss in more detail some strengths and weaknesses of current methods that seek to improve FF efficiency. Then, we formalize the problem we are trying to solve, along with its motivation in Section 3. We present our novel approach in Section 4.2, which requires a thorough examination of the surprising phenomenon of flocking shared by many LLMs in Section 4.1. Our rigorous experiments demonstrate GRIFFIN preserves performance on classification and generation even after removing 50% of FF neurons (Section 5.1), all while having lower latency (Section 5.2). For instance, GRIFFIN reduces the number of active parameters in Llama 2 13B from 13B to 8.8B during generation to improve latency by 1.25$\times$ with almost no loss in performance. Finally, we show our method’s incredible scalability and robustness in several ablation studies (Section 5.3).
@@ 2 Background
The observation that transformer FF blocks produce sparse activations is not new [GSBL20, DLBZ22, LYB+22, DCC23, LWD+23]. In ReLU-based LLMs like OPT [ZRG+22], the activations can be exceptionally sparse and become more apparent for larger models [LWD+23]. As more models use non-sparse activation functions like GLU variants [Sha20], it is difficult for neurons to have no contribution to the output since these functions do not have an interval that maps to zero. Without exact sparsity, the efficacy of these methods becomes limited. As such, this has ushered a wave of models that are either adapted from available models [ZLL+21, LLLC21, MAM+23, ZBC+24, JSR+24] or trained from scratch [FZS22] which can produce activations that are exactly zero with little to no performance loss. Even so, these methods require considerable amounts of computational resources.
@@
Pruning [LDS89] is one sparsity-guided way to tackle compute and memory bottlenecks of models. Previously, the common method would be some variation of iteratively rounding weights down to zero based on some score and retraining to recover any lost performance [FC18, BGOFG20, LGW+21, LZH+24]. While this can result in most parameters being pruned, this method comes with a few issues. First, with the increasing scale of LLMs, retraining becomes impractical for most. Fortunately, cheap methods to effectively prune LLMs have been developed [FA23, SLBK23, JLC+23, DKK+24]. The second issue is that unless pruning is done in a structured manner [XZC22, SWSL23, MFW23, LYZ+23, XGZC23], it is difficult to see real computational savings, yet structured pruning often leads to much more severe performance degradation. Third, pruning usually enforces sparsity to be static at inference which can be strong assumption since FF blocks are widely believed to contain the model’s memory [GSBL20].
@@
Making sparsity more dynamic has motivated the design of mixture-of-experts (MoEs) [JJNH91] to avoid computing low-impact features in trained models at varying granularities [ZLL+21, LWD+23, PSBW23, CIS23, YYB+24, ZBC+24]. The main idea of MoEs is to use a gating function to identify a small subset of neurons that will be used to compute the layer’s output, an adaptive form of pruning. In the ideal case, all active neurons are selected and inactive neurons are ignored for each input. However, current methods either require training or rely on ReLU or ReLU-like activation functions.
@@ 3 Problem Formulation
This section contains an overview of different components of the FF block followed by a formulation of the FF compression problem which our method aims to tackle. Since FF blocks operate identically and independently for each token unlike attention, we begin with defining the FF block with a single column vector input $\bm{x}\in\mathbb{R}^{D}$:
@@
where $\odot$ signifies element-wise multiplication. For all examples, $\bm{W}_{1},\bm{W}_{g}\in\mathbb{R}^{D_{\text{FF}}\times D}$ and $\bm{W}_{2}\in\mathbb{R}^{D\times D_{\text{FF}}}$ where typically, $D_{\text{FF}}\gg D$. We refer to $\bm{z}=\text{FF}_{1}(\bm{x})$ as the FF activations.
A popular method to compress FF blocks (which is also the goal of GRIFFIN, adaptive neuron pruning, and many MoE methods) is to find $\widehat{\bm{W}}_{1}\in\mathbb{R}^{k\times D}$, $\widehat{\bm{b}}_{1}\in\mathbb{R}^{k}$, and $\widehat{\bm{W}}_{2}\in\mathbb{R}^{D\times k}$ (additionally $\widehat{\bm{W}}_{g}\in\mathbb{R}^{k\times D}$ and $\widehat{\bm{b}}_{g}\in\mathbb{R}^{k}$ if needed) where $k<D_{\text{FF}}$ such that when the FF block is reparameterized with these matrices, the output value is preserved. Namely, for
@@
$\text{FF}(\bm{x})\approx\widehat{\text{FF}}_{2}(\widehat{\text{FF}}_{1}(\bm{x}))$, and similarly for FF blocks with non-GLU functions. For instance, in the MoE setting, these smaller matrices can vary from token to token and are actually selections of chunks of the original structures. Crucially, a solution to this problem leads to multiplication with smaller matrices which are naturally more efficient on GPUs and TPUs [FSH04, WWB19]. Note that for all equations defined in this section, they operate independently on each row of a length $S$ sequence input $\bm{X}\in\mathbb{R}^{S\times D}$ (e.g. the activations for a sequence are $\bm{Z}=\text{FF}_{1}(\bm{X})\in\mathbb{R}^{S\times D_{\text{FF}}}$).
@@ 4.1 Observing Flocking
Flocking arises when we look at the relative impact of each neuron per token within a sequence. To see this, we normalize rows of $\bm{Z}$ to be unit vectors to construct $\overline{\bm{Z}}\in\mathbb{R}^{S\times D_{\text{FF}}}$ (i.e. $[\overline{\bm{Z}}]_{i}=[\bm{Z}]_{i}/\|[\bm{Z}]_{i}\|_{2}$), the relative activations. We show example relative activation magnitudes for a sequence in Llama 2 7B and Gemma 7B in Figure 1. Since there are distinct vertical streaks, this intriguingly implies that activations that have relatively greater weight are common across all tokens in a sequence. Notably, Llama 2 7B and Gemma 7B use SwiGLU and GEGLU activations, respectively, along with other major architecture differences. We call this phenomenon flocking, like highly organized groups of birds, and we observe this in virtually all FF layers (see Appendix E) and randomized inputs (see Appendix C).
While relative activations magnitudes are shared within a sequence, they are not generally shared between sequences. We show this by taking the $\ell_{2}$-norm of $\overline{\bm{Z}}$ along the token axis to obtain a length $D_{\text{FF}}$ vector for each sample or sequence, roughly capturing the contribution of each FF neuron throughout a sequence. Taking the top-$k$ of this for each sample at each layer, we find the Jaccard similarity between two sequences based on the indices selected for different $k$. In other words, we compute the intersection over union of every unique pair of top-$k$ sets. Higher values indicate more similar top-$k$ sets. From Figure 2 where we aggregate Jaccard similarities across WikiText [MXBS16] samples, we observe a lack of inter-sample activation similarities for the vast majority of layers in Llama 2 7B and Gemma 7B, unless the sets of selected neurons are large. This lack of consistency implies statically pruning entire FF neurons without retraining would be less effective than a more adaptive method.
@@ 4.2 GRIFFIN Algorithm
Using our insight on flocking, we introduce GRIFFIN as a simple general purpose and training-free adaptive structured pruning method for efficient generation, captured in Figure 3. In a nutshell, we select neurons during the prompt phase of each sample which are then used for the entire duration of the generation phase. This effective approach is based on a key observation on flocking: since tokens within a sequence share activation patterns, the prompt and generated tokens will also share activation patterns.
Prompt Phase Expert Neuron Selection.
Our expert neurons are chosen at the sequence level, so we need to consider the dynamics of the entire input sequence rather than just a single token when choosing our neurons. To select expert neurons, we need a statistic $\bm{s}\in\mathbb{R}^{D_{\text{FF}}}$ to inform us of the importance of each neuron. At the prompt phase, we do this by taking the $\ell_{2}$-norm of $\overline{\bm{Z}}$ along the token axis:
@@
Taking the indices of the top-$k$ across $\bm{s}$ gives us the neurons we will use for this sample’s generation phase which make up the set $\mathcal{E}$. Using the expert neurons in $\mathcal{E}$, we can find $\widehat{\bm{W}}_{1},\widehat{\bm{b}}_{1},\widehat{\bm{W}}_{g},\widehat{\bm{b}}_{g},\text{ and }\widehat{\bm{W}}_{2}$ by selecting corresponding rows and columns in $\bm{W}_{1},\bm{b}_{1},\bm{W}_{g},\bm{b}_{g}$, and $\bm{W}_{2}$, respectively. This is done for every FF block during the prompt phase. Illustrated in Appendix A, $\bm{s}$ highlights neurons consistently activated at relatively high intensities.
Generation with Expert Neurons.